OpenStudy (anonymous):

Find the derivative. f(x)=x^2(x^3+3x^2) I got 120x+72. Anyone care to check my answer?

OpenStudy (lgbasallote):

did you use product rule?

OpenStudy (anonymous):

yea.

OpenStudy (turingtest):

why can't we just distribute first?

OpenStudy (lgbasallote):

i got something different

OpenStudy (lgbasallote):

oh lol of course

OpenStudy (anonymous):

What did you get Igbasallote?

OpenStudy (lgbasallote):

just distribute...it's faster...youll get same answer anyway... \[\frac{d}{dx} (x^5) + \frac{d}{dx} (3x^2)\]

OpenStudy (anonymous):

I did. I know how to do it, i just want to know where i did the mistake lol o_O

OpenStudy (lgbasallote):

but if you prefer product rule... \[x^2 \frac{d}{dx} (x^3 + 3x^2) + (x^3 + 3x^2)\frac{d}{dx} (x^2)\]

OpenStudy (lgbasallote):

derivative of x^5 is 5x^4 due to the power rule \(\frac{d}{dx}x^n = nx^{n-1}\)

OpenStudy (anonymous):

continueeeee. I got 20x^3 after that.

OpenStudy (lgbasallote):

btw that\s supposed to be \(\frac{d}{dx} (3x^4)\) not 3x^2

OpenStudy (lgbasallote):

i mistyped

OpenStudy (lgbasallote):

derivative of 3x^4 = \((3 \times 4)x^{4-1}\)

OpenStudy (anonymous):

yea after that I got 5x^4+12x^3

OpenStudy (lgbasallote):

^that's right

OpenStudy (anonymous):

then 20x^3+36x^2

OpenStudy (anonymous):

hello?

OpenStudy (lgbasallote):

why

OpenStudy (lgbasallote):

why did you make it 20x^3 + 36x^2?

OpenStudy (anonymous):

because the exponent times 5 and then decrease the exponent by 1. So 5*4=20, then x^4-1=20x^3

OpenStudy (anonymous):

and i did the same for the next number.

OpenStudy (lgbasallote):

are you looking for second derivative?

OpenStudy (anonymous):

I'm looking for the most simplified derivative possible.

OpenStudy (lgbasallote):

that's as simplified you can get...if you take the derivative it's no longer equal

OpenStudy (lgbasallote):

5x^4 + 12x^3 is enough

OpenStudy (anonymous):

But I'm required to go further to find it until the 5th derivative. This is only the 2nd or 3rd.

OpenStudy (lgbasallote):

oh..until 5th derivative? well 2nd derivative = 20x^3 + 36x^2 3rd derivative = 60x^2 + 72x 4th derivative = 120 x + 72 5th = 120

OpenStudy (anonymous):

Yup I got that for my 4th lol. Thanks. That's all I needed to know lol.